# Human-inspired Perspectives: A Survey on AI Long-term Memory

### v1

@@ 3.1.3 Long-term Memory
    Episodic Memory. Tulving et al. [tulving1972episodic] defined episodic memory in 1972 as a type of memory specific to individuals, related to time and space, involving personal experiences, events, and contexts. In 1973, Tulving et al. [tulving1973encoding] noted that input stimuli such as sound and visual information are processed for episodic memory storage through specific encoding, which combines perceived information with specific contexts and experiences to form memory trace (discussed in the leading paragraph of Sec. 3.1.3) that aids in subsequent retrieval. For successful recall of this trace, similar input stimuli need to be provided. For example, adults often recall childhood experiences upon seeing childhood photos, which stimulates the recall of relevant long-term memory traces. In 1985, Tulving et al. [tulving1985memory] explored the relationship between memory and consciousness, suggesting that episodic memory is associated with autonoetic consciousness. This form of consciousness allows individuals to associate the present situation with specific personal events from the past, with an awareness that the situation are genuine parts of their past experiences. Furthermore, episodic memory also facilitates learning from past experiences and applying this knowledge to future decisions, thereby enhancing an individual’s ability to adapt to their environment [tulving2002episodic].
@@
    Semantic Memory. According to Tulving et al.’s perspective [tulving1972episodic] in 1972, semantic memory encompasses an individual’s knowledge about words and other linguistic symbols, including the relationships between them, and the rules, formulas, and algorithms for manipulating them. It functions to receive, retain, and convey semantic information about words and concepts. Unlike episodic memory, which is tied to personal experiences, semantic memory deals with general information. An example of semantic memory is the long-term memory formed by learning knowledge in a specific field or by extracting general rules from past experiences. In 1985, Tulving et al. [tulving1985memory] further explored the relationship between memory and consciousness, indicating that the type of consciousness associated with semantic memory is termed noetic consciousness. Noetic consciousness is an internal, non-physical state of awareness that allows individuals to perceive certain objects without external stimuli and is more closely associated with intuition, insight, and internal knowledge. For example, when a person recalls the definition of a word or a mathematical formula, they do so through noetic consciousness, which does not need to be called up based on a certain scene. This contrasts with autonoetic consciousness, which connects the present scene to episodic memory.
@@
    Procedural Memory. Tulving et al. [tulving1985memory] describe procedural memory as the memory developed from learning motor skills through feedback and perceptual abilities. This type of memory is intrinsically linked to the immediate temporal and spatial contexts and does not depend on information beyond the current stimuli. This allows for instantaneous reactions based on direct perception. For example, learning to ride a bicycle involves the sensation of losing balance, which triggers discomfort and gradually leads to the development of balancing skills. Once these skills are fully acquired, individuals can effortlessly maintain balance while riding, demonstrating the role of procedural memory in learning motor skills.
@@ 3.2 Human Memory Processing
However, as mentioned earlier, memory storage and retrieval only address how memory are formed and utilized. There also exists a mechanism facilitating memory extinction, which we refer to as Memory Forgetting. Specifically, the research by Ebbinghaus et al. [ebbinghaus1885memory] in the late 19th century sparked continued interest among subsequent neuroscientists in the field of memory forgetting, leading to a series of experimental analyses and ongoing attention [jost1897law, wixted1997genuine, anderson1997memory].
From a neuroscience perspective, we observe more fine-grained mechanisms of memory processing. Several studies highlight the critical role of the hippocampus in this process. Lavenex et al. [lavenex2000hippocampal] revealed in 2000 that interactions between the hippocampus and the neocortex are essential for storing long-term declarative memory, which includes both episodic and semantic memory (discussed in Sec. 3.1.3). However, there is debate about whether the hippocampus is required for retrieving all episodic memory or only for the creation of new episodic memory [wiltgen2010hippocampus]. In 2004, Deisseroth et al. [deisseroth2004excitation] demonstrated that excitatory stimuli can increase the proportion of neural stem cells/progenitor cells (NPCs) in the adult hippocampus differentiating into neurons. These new neurons may participate in the storage of new memory, and their addition may accompany the replacement of synaptic connections of old neurons, thereby accelerating the forgetting of old memory. Furthermore, Rolls et al. [rolls2006spatial] in 2006 indicated that spatial view cells in the hippocampus are involved in the storage and retrieval of spatial location-related memory. More recently, in 2018, Meira et al. [meira2018hippocampal] found that activity in a specific region of the hippocampus, the dorsal CA2, is crucial for the encoding and recall of social memory, which pertains to an individual’s ability to recognize and remember social interactions and partners. On the other hand, Weinberger et al. [weinberger2004specific] demonstrated in 2004 that the primary auditory cortex can acquire and retain specific memory traces about the behavioral significance of certain sounds, the plasticity changes in the cortex strengthens over time and specific memory-related sound stimuli can activate particular neural networks within the primary auditory cortex. This highlights the cortex’s crucial role in the storage, consolidation, and retrieval of long-term memory.
@@ 3.2.1 Memory Storage
Based on the perspective of Tulving et al. [tulving1973encoding], memory storage is achieved by processing input stimuli through specific encoding (detailed in the episodic part of Sec. 3.1.3). According to Atkinson et al. [shiffrin1969memory], sensory register, short-term store, and long-term store have distinct methods for storing and encoding information at different stages of the memory process. The long-term store, in particular, is viewed as a permanent repository of information. It encodes memory through four distinct encoding strategies: “meaning”, “association”, “repetition”, and “organization”, as discussed below:
@@ 3.2.2 Memory Retrieval
According to Tulving et al. [tulving1973encoding], memory retrieval involves a conscious search and identification of information stored in memory. They emphasized that this process is not merely the activation of learned associations or the recall of stored traces. Instead, it involves complex interactions between stored information and specific features of the current retrieval context. Providing cues similar to the original stimuli that formed the memory can enhance retrieval effectiveness. An example of this is individuals visiting a place they have previously stayed at, which is likely to evoke the events they experienced there.
@@ 4.2.2 Retrieval of Parametric Memory
Figure 5: The similarities between different types of AI long-term memory and human long-term memory. Types of Non-Parametric Memory refer to Section 4.1.1, and types of Parametric Memory refer to Section 4.2.3.
@@ 6 A New Cognitive Architecture for Long-term Memory
    Adaptive Forgetting. The long-term memory forgetting process in SALM is depicted in Part C of Fig. 8. In non-parametric memory, the forgetting adapter enhances the effectiveness of active forgetting to improve essential information retrieval (Sec. 4.1.3). This enhancement can be achieved by refining the selection of forgetting targets based on specific objectives. For parametric memory, forgetting adapters primarily prevent catastrophic forgetting [aleixo2023catastrophic] through various methods. These may include optimizing the selection of rehearsal samples [lopez2017gradient, nguyen2018variational, de2019episodic, prabhu2020gdumb, roscow2021learning] and identifying high-importance parameter subsets [Adel2020Continual, xue2022meta].
@@ 7.2 Application of AI Long-term Memory
    Video Understanding. Video understanding refers to tasks related to video analysis such as recognition of entities and actions within videos and video content summarization [jiao2021new, wu2021towards, apostolidis2021video, ju2022prompting]. This field has broad applications, such as in movie and surveillance analysis, autonomous driving, and aerospace remote sensing systems. Incorporating long-term memory significantly enhances video understanding by enabling the storage and retrieval of key video information for subsequent tasks [cheng2022xmem, shen2023encode]. A typical video understanding pipeline driven by long-term memory converts video frames into feature vectors, which can later be retrieved to support various tasks [cheng2022xmem, mezghan2022memory, shen2023encode, song2023moviechat, he2024ma]. These frames contain situational information from scenes, which is regarded as episodic memory. Moreover, some video understanding systems improve their performance by integrating external information sources (semantic memory), such as video clip captions [yang2023care] and related videos [jilan2024egoInstructor]. Our proposed SALM offers a promising approach for developing video understanding systems capable of managing various types of AI long-term memory. The performance of SALM-based systems can be evaluated using various datasets related to video understanding. For instance, the Ego4D dataset [grauman2022ego4d] captures daily activities from a first-person perspective, supporting tasks such as indexing past experiences, analyzing present interactions, and anticipating future activities. The Ego-Exo4D dataset [grauman2024ego], which includes both egocentric and exocentric views of skilled human activities, is designed for tasks like activity understanding, proficiency estimation, cross-view translation, and 3D hand/body pose estimation. The Replica dataset [straub2019replica] is commonly used for tasks like 2D/3D semantic segmentation and geometric inference in indoor spaces. Additionally, tools like Project Aria [engel2023project] contribute to dataset creation by enabling egocentric multi-modal data recording and streaming.

### v2

@@ 3.1.3 Long-term Memory
    Episodic Memory. Tulving et al. [tulving1972episodic] defined episodic memory in 1972 as a type of memory specific to individuals, related to time and space, involving personal experiences, events, and contexts. In 1973, Tulving et al. [tulving1973encoding] noted that input stimuli such as sound and visual information are processed for episodic memory storage through specific encoding, which combines perceived information with specific contexts and experiences to form memory trace (discussed in the leading paragraph of Sec. 3.1.3) that aids in subsequent retrieval. For successful recall of this trace, similar input stimuli need to be provided. For example, adults often recall childhood experiences upon seeing childhood photos, which stimulates the recall of relevant long-term memory traces. Furthermore, episodic memory also facilitates learning from past experiences and applying this knowledge to future decisions, thereby enhancing an individual’s ability to adapt to their environment [tulving2002episodic]. The hippocampus in the brain plays a vital role in learning episodic memory by encoding them through a process that separates similar personal experiences and stores them distinctly, thereby minimizing interference between different episodic memory [marr1991simple, o1994hippocampal, o2014complementary].
@@
    Semantic Memory. According to Tulving et al.’s perspective [tulving1972episodic] in 1972, semantic memory encompasses an individual’s knowledge about words and other linguistic symbols, including the relationships between them, and the rules, formulas, and algorithms for manipulating them. It functions to receive, retain, and convey semantic information about words and concepts. Unlike episodic memory, which is tied to personal experiences, semantic memory deals with general information. An example of semantic memory is the long-term memory formed by learning knowledge in a specific field or by extracting general rules from past experiences. The management of semantic memory involves multiple brain regions, such as the left prefrontal cortex (LPC) being involved in the retrieval of semantic information, and the temporal lobes storing specific information related to objects [hodges1992semantic, baldo1998letter, martin2001semantic].
@@
    Procedural Memory. Tulving et al. [tulving1985memory] describe procedural memory as the memory developed from learning motor skills through feedback and perceptual abilities. This type of memory is intrinsically linked to the immediate temporal and spatial contexts and does not depend on information beyond the current stimuli. This allows for instantaneous reactions based on direct perception. For example, learning to ride a bicycle involves the sensation of losing balance, which triggers discomfort and gradually leads to the development of balancing skills. Once these skills are fully acquired, individuals can effortlessly maintain balance while riding, demonstrating the role of procedural memory in learning motor skills. Brain regions such as the supplementary motor area (SMA), primary motor cortex (M1), prefrontal cortex, and cerebellum play crucial roles in skill consolidation for the formation of procedural memory [censor2014cortico].
@@ 3.2 Human Memory Processing
Beyond storage and retrieval, forgetting also plays a vital role in the understanding of memory. Specifically, the research by Ebbinghaus et al. [ebbinghaus1885memory] in the late 19th century sparked continued interest among subsequent neuroscientists in the field of memory forgetting, leading to a series of experimental analyses and ongoing attention [jost1897law, wixted1997genuine, anderson1997memory].
From a neuroscience perspective, experimental analyses of cellular mechanisms in memory processing continue to uncover the anatomical and physiological foundations of memory. Many studies highlight the critical roles of the hippocampus, the neocortex, and their interaction in supporting long-term memory [marr1991simple, lavenex2000hippocampal, martin2001semantic, ji2007coordinated, o2014complementary, censor2014cortico] (Discussed in Sec. 3.1.3 and the subsequent subsections). In brains, memory stability necessarily reflects the lifecycle of neurons. For example, Deisseroth et al. [deisseroth2004excitation] demonstrated that excitatory stimuli can increase the proportion of neural stem cells/progenitor cells (NPCs) in the adult hippocampus differentiating into neurons. The genesis of new neurons may participate in the storage of new memory, and their addition may accompany the replacement of synaptic connections of old neurons, thereby accelerating the forgetting of old memory. Meanwhile, different subfields of the hippocampus have been implicated in supporting different memory modalities. For example, dorsal CA2 of the mouse hippocampus may preferentially contribute to the encoding and recall of social memories related to the identity of conspecifics [meira2018hippocampal]. Furthermore, Weinberger et al. [weinberger2004specific] demonstrated in 2004 that the primary auditory cortex can acquire and retain specific memory traces about the behavioral significance of certain sounds, the plasticity changes in the cortex strengthens over time and specific memory-related sound stimuli can activate particular neural networks within the primary auditory cortex. This highlights the cortex’s crucial roles in the storage, consolidation, and retrieval of long-term memory.
@@ 3.2.1 Memory Storage
Based on the perspective of Tulving et al. [tulving1973encoding], memory storage is achieved by processing input stimuli through specific encoding (detailed in the episodic part of Sec. 3.1.3). The hippocampus and cortex of the brain are involved in the encoding of memory, with the hippocampus encoding unique episodic information through its pattern separation properties and the cortex encoding more generalized information through its distributed representations [marr1991simple, o1994hippocampal, winocur2010memory, o2014complementary]. Sensory register (Sec. 3.1.1), short-term store (Sec. 3.1.2), and long-term store (Sec. 3.1.3) have distinct methods for storing and encoding information at different stages of the memory process [shiffrin1969memory]. The long-term store, in particular, encodes memory through four distinct encoding strategies: “meaning”, “association”, “repetition”, and “organization”, as discussed below:
@@ 3.2.2 Memory Retrieval
According to Tulving et al. [tulving1973encoding], memory retrieval involves a conscious search and identification of information stored in memory. They emphasized that this process is not merely the activation of learned associations or the recall of stored traces. Instead, it involves complex interactions between stored information and specific features of the current retrieval context. Providing cues similar to the original stimuli that formed the memory can enhance retrieval effectiveness. An example of this is individuals visiting a place they have previously stayed at, which is likely to evoke the events they experienced there. The hippocampus of brain is involved in the complex process of retrieving specific memory based on cues [marr1991simple, o2014complementary].
@@ 4.2.3 Forgetting of Parametric Memory
Figure 5: The similarities between different types of AI long-term memory and human long-term memory. Types of Non-Parametric Memory refer to Sec. 4.1.1, and types of Parametric Memory refer to Sec. 4.2.3.
@@ 6 A New Cognitive Architecture for Long-term Memory
    Adaptive Forgetting. The long-term memory forgetting process in SALM is depicted in Part C of Fig. 8. In non-parametric memory, the forgetting adapter enhances the effectiveness of active forgetting to improve essential information retrieval (Sec. 4.1.3). This enhancement can be achieved by refining the selection of forgetting targets based on specific objectives. For parametric memory, forgetting adapter primarily prevents catastrophic forgetting [aleixo2023catastrophic] through various methods. These may include optimizing the selection of rehearsal samples [lopez2017gradient, nguyen2018variational, de2019episodic, prabhu2020gdumb, roscow2021learning] and identifying high-importance parameter subsets [Adel2020Continual, xue2022meta].
@@ 7.2 Application of AI Long-term Memory
    Video Understanding. Video understanding refers to tasks related to video analysis such as recognition of entities and actions within videos and video content summarization [jiao2021new, wu2021towards, apostolidis2021video, ju2022prompting]. This field has broad applications, such as in movie and surveillance analysis, autonomous driving, and aerospace remote sensing systems. Incorporating long-term memory significantly enhances video understanding by enabling the storage and retrieval of key video information for subsequent tasks [cheng2022xmem, shen2023encode]. A typical video understanding pipeline driven by long-term memory converts video frames into feature vectors, which can later be retrieved to support various tasks [cheng2022xmem, mezghan2022memory, shen2023encode, song2023moviechat, he2024ma]. These frames contain situational information from scenes, which can be regarded as episodic memory. Moreover, some video understanding systems improve their performance by integrating external information sources (semantic memory), such as video clip captions [yang2023care] and related videos [jilan2024egoInstructor]. Our proposed SALM offers a promising approach for developing video understanding systems capable of managing various types of AI long-term memory. The performance of SALM-based systems can be evaluated using various datasets related to video understanding. For instance, the Ego4D dataset [grauman2022ego4d] captures daily activities from a first-person perspective, supporting tasks such as indexing past experiences, analyzing present interactions, and anticipating future activities. The Ego-Exo4D dataset [grauman2024ego], which includes both egocentric and exocentric views of skilled human activities, is designed for tasks like activity understanding, proficiency estimation, cross-view translation, and 3D hand/body pose estimation. The Replica dataset [straub2019replica] is commonly used for tasks like 2D/3D semantic segmentation and geometric inference in indoor spaces. Additionally, tools like Project Aria [engel2023project] contribute to the development of datasets by enabling the recording and streaming of egocentric multi-modal data.